OpenStudy (anonymous):

Josh and his cousin Lucy are picking apples in their grandparents' orchard. Josh has filled 13 baskets with apples and is filling them at a rate of 3 baskets per hour. Lucy has 9 full baskets and will continue picking at 5 baskets per hour. Once the cousins get to the point where they have filled the same number of baskets, they will carry them to the barn and then eat lunch. Which equation is part of the system of equations? A.3x + 5y = 13 B.3x + 13 = y C.9x - 3y = 13 D. 5x + 13 = y

OpenStudy (anonymous):

OK, if you pull out the numbers with the proper variables, what do you think?

OpenStudy (anonymous):

So they are both trying to fill 13 baskets?

OpenStudy (anonymous):

@C0stell0

OpenStudy (anonymous):

Josh J Lucy L Hours h Josh's rate is 3, but he has 13 already, so: 3h+13=J Lucy, on the other hand is 9 already and 5 an hour: 5h+9=L When they will have the same is when J = L, so: 3h+13=5h+9 So I am not sure why they are complicating this with extra variables.